Below is a class distribution table for the weights(in kg) of athletes who registered for a wrestling competition.



Weight Frequency



70-74 4



75-79 7



80-84 11



85-89 8





Find the mean weight of the wrestlers who will be competing in the tournament. Write your complete solution below



1
Expert's answer
2022-04-13T15:33:16-0400

"mu=\frac{72*4+77*7+82*11+87*8}{4+7+11+8}=\frac{2425}{30}=" 80.33.